TUTTI 1. Find the distance betweenthe two planes. x – 2y + 3z = -4 and x – 2y + 3z = 2.
Answer #1

. Given plancs - x-2y+32=-4 - 0 x-2y+32=2 If Ax+By+C2+120 and Axtøy +CZ +D250 is a plane equation, then distance between plan* The they given plancs satisfied are the paralele plones since condition, that is are Two plones; Axt a x+azy + Az Z=0 b,x+

Thanks

Please do rate and review. Keep HomeworkLibing.